If we have $4000 to put in the bank earning 4.5 % interest compounded 2 times per year; how long will it take before we double our money?
Your money would double in years.
If we have $P o to put in the bank earning 4.5 % interest compounded 2 times per year; how long will it take before we double our money?
Your money would double in years

Answer by lwsshak3(11628)

If we have $4000 to put in the bank earning 4.5 % interest compounded 2 times per year; how long will it take before we double our money?
Your money would double in years.
If we have $P o to put in the bank earning 4.5 % interest compounded 2 times per year; how long will it take before we double our money?
Your money would double in years
Compound interest formula: A=P(1+r/n)^nt,P=initial investment, r=interest rate, n=number of compounding periods per year, A=amt after t-yrs
First problem:
P=4000
r=.045
n=2
A/P=2=(1+.045/2)^2t
2=(1.0225)^2t
log 2=2tlog(1.0225)
take log of both sides
2t=log2/log(1.0225)=31.15
t=15.6 yrs
how long will it take before we double our money? 15.6 yrs
2nd problem: same answer as first problem.(note:amt of initial investment does not matter)